1)    A car accelerates from 2.0 m/s to speed of 10.0 m/s in 4.0 sec. If the rate of acceleration is constant, determine:

a)    Acceleration

b)    Distance traveled



1
Expert's answer
2021-05-03T10:34:22-0400

The acceleration is the rate of change of speed:


"a=\frac{v_f-v_i}{t}=\frac{10-2}{4}=2\text{ m/s}^2."

The distance traveled is


"d=\frac{v_f^2-v_i^2}{2a}=24\text{ m}."